OpenStudy (anonymous):

what is an equation in slope-intercept form for the line that passes through the points (1,-3) and (3,1)? I got y = 2x + 5

OpenStudy (anonymous):

The 5 should be negative

OpenStudy (anonymous):

That's what I thought, but when I looked at the answer choices, none of them were negative 5's. These were the choices: y = 3x + 1 y = x – 3 y = 2x + 5 y = 2x – 5

OpenStudy (anonymous):

The last one is right

OpenStudy (anonymous):

Okay, I thought since the slope intercept form equation was y = mx + b, that it had to be a addition sign.

OpenStudy (anonymous):

Thats just the form you need to have, the b could be positive or negative it just depends on what line you're doing

OpenStudy (anonymous):

Thanks.